OpenStudy (anonymous):

given the function f(x)=4(2)^x, section A is from x=1 to x=2 and section B is from x=3 to x=4 Part A: Find the Average Rate of change of each section Part B: How many times greater is the average rate of change of section B than section A? Explain Why one rate of change is greater than the other.

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@Agl202

OpenStudy (anonymous):

@AaronAndyson

OpenStudy (anonymous):

@aloud

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@partyqueen

OpenStudy (anonymous):

@paulalovee

OpenStudy (anonymous):

@Wendy.Ivette11714

OpenStudy (anonymous):

@ganeshie8 can you help this person plz im not sure how to explane or do this problem right?

OpenStudy (anonymous):

@rockstar0765

OpenStudy (anonymous):

@Ashleyrh1999

OpenStudy (anonymous):

can anyone help?

OpenStudy (anonymous):

@dr0zier99

OpenStudy (anonymous):

@ConnexusStudent0912

OpenStudy (dr0zier99):

are you trying to evaluate it

OpenStudy (dr0zier99):

Substitute the value of the variable into the equation and simplify f(3)=64

OpenStudy (dr0zier99):

section a: f(x)=4(2)^x 1.......8 2.......16 section b: 3.......32 4.......64

OpenStudy (dr0zier99):

section b is 4 times greater than section a

OpenStudy (anonymous):

if i simplify f(3)=64 I get 0

OpenStudy (michele_laino):

I think that the definition of average rate \(r\) of change, can be this: \[\Large r = \frac{{f\left( {{x_2}} \right) - f\left( {{x_1}} \right)}}{{{x_2} - {x_1}}}\] where: \[\Large f\left( x \right) = 4 \cdot {2^x}\]

OpenStudy (anonymous):

Michele _Laino is that for Part A?

OpenStudy (michele_laino):

it is for parts A, and B for example, for part A, we can write: \(x_1=1,\,x_2=2\)

OpenStudy (anonymous):

oh ok

OpenStudy (michele_laino):

whereas for part B, we can use the same formula, with \(x_1=3,\,x_2=4\)

OpenStudy (dr0zier99):

so i was correct when i said that part b was 4 times greater than part a

OpenStudy (anonymous):

what about the last part for part b explain why one rate of change is greater than the other?

OpenStudy (anonymous):

Yes you were correct when you said that

OpenStudy (anonymous):

Thanks for helping guys but I don't know who to give the medal to

OpenStudy (dr0zier99):

no problem

OpenStudy (michele_laino):

please give it to @dr0zier99 he/she worked more than me :)

OpenStudy (dr0zier99):

ima boy

OpenStudy (dr0zier99):

XD

OpenStudy (michele_laino):

ok! he worked more than me!

OpenStudy (dr0zier99):

still i give medel to u cause u still helped